### v3

# NeuronMM: High-Performance Matrix Multiplication for LLM Inference on AWS Trainium

Abstract.
[…]
submissionid: 219
1. Introduction
Large Language Models (LLMs) have achieved remarkable success across a wide range of text-based tasks (liu2024deepseek; yang2025qwen3). Yet, the steady growth in their parameter counts and architectural complexity makes deployment increasingly prohibitive, particularly in resource-constrained environments. These challenges have driven extensive research into model compression (agarwal2024policy; frantar2023sparsegpt; dettmers2022gpt3; dettmers20218; sanh2020movement; zhang2020accelerating; li2023symbolic; chen2022disco) and efficient hardware design (wan2023efficient; zhu2024survey; jouppi2020domain; bai2023transformer; tzanos2022hardware; tuli2023acceltran; liu2023hardsea; ji2024beta; tuli2023acceltran).
AI accelerators (jouppi2023tpu; coburn2025meta; firoozshahian2023mtia), customized to AI workloads, provide cost-effective and high-performance solutions for training and inference. Trainium is an AI accelerator recently developed by Amazon Web Services (AWS). It has been reported that Trainium can deliver 30–40% lower cost while providing performance comparable to GPU-based EC2 instances (aws_trainium). Each Trainium2 device (the most advanced Trainium) integrates two NeuronCores, each delivering up to 95 TFLOPS of FP16/BF16 compute capability, comparable to NVIDIA A100 GPUs at roughly 60% of the cost. Such a cost–effective advantage makes Trainium an attractive platform for LLM training and inference (fu2024distributed; fan2024hlat). Furthermore, Trainium, as a typical systolic-array architecture, features a programmable memory hierarchy (including two types of on-chip SRAMs and off-chip HBM). It also provides a rich set of specialized compute engines tailored for various AI operators. Such hardware heterogeneity gives programmers a lot of flexibility to explore for better performance.
However, leveraging Trainium architecture for high performance can be challenging. First, as a systolic array architecture, Trainium must repeatedly go through a load-compute-store cycle to accommodate the small capacity of its on-chip SRAM. This design causes frequent data movement across the memory hierarchy, whose overhead can often be larger than the computation time spent in various compute engines in Trainium. In addition, allocating too much data to the on-chip SRAM can lead to implicit “memory spill” to HBM, which stalls the compute engines. On the other hand, underutilizing the on-chip SRAM wastes its high memory bandwidth and lowers the overall system throughput. Second, the programmer must carefully align a tensor’s logical shape with Trainium’s physical memory layout. Misalignment often requires tensor transposes, which incur costly data transfers between HBM and on-chip SRAM.
In this paper, we design high-performance matrix multiplication (matmul), a dominating compute kernel in LLM, for LLM inference on Trainium. We call our system, NeuronMM. Building high-performance matmul for LLM inference on Trainium, we face the challenges discussed above. To reduce data movement overhead, we apply Singular Value Decomposition (SVD) to the weight matrices in LLM. By factoring a large weight matrix into two but retaining the top singular values and their corresponding singular vectors, we obtain a low-rank approximation to the original weight matrix but with smaller matrix size, hence leading to reductions of data movement. To make the application of SVD aligned with the capacity of on-chip SRAM in Trainium and maximize SRAM utilization, we employ a three-level hierarchical data layout (i.e, tile, block, and strip) and introduce a block-wise SVD.
After applying SVD to the weight matrix $W$ ($W\approx UV$), the large matmul in LLM ($XW$, where $X$ is input embedding) is transformed to $XUV$, a sequence of two matmuls. Naively implementing it on Trainium suffers from the overhead of the data movement and data layout transpose. In particular, materializing the intermediate result (the output of the first matmul in the sequence) causes data reloading from HBM because of small SRAM capacity and the necessity of storing the output of the first matmul in HBM. To avoid this problem, we can fuse the two matmuls in the sequence without the materialization of the intermediate result; this approach recomputes the blocks of the intermediate result when needed without storing them in HBM. However, it brings a new I/O overhead of loading sources blocks of $X$ and $U$ to compute each block of the intermediate result, which outweighs the benefit of recomputation. Furthermore, the transpose must happen between the two matmuls in the sequence, leading to extra overhead.
To address the above problem, NeuronMM introduces a new kernel fusion method for Trainium, named TrainiumFusion. It is featured with three major techniques. First, TrainiumFusion introduces an SRAM-capacity-aware caching strategy to eliminate recomputation penalty. This strategy caches multiple rows of the intermediate matrix on the SRAM based on its capacity, and carefully reuses it when generating the output blocks with the corresponding column strips in the source blocks. This method avoids recomputation and frequent data movement. Second, TrainiumFusion reduces matrix transpose by leveraging the matrix-identify property without impacting the result correctness. Third, TrainiumFusion computes the matmul sequence by blocks in combination with its caching strategy and DMA-assisted result accumulation in SRAM. We further develop performance modeling to quantify the relationship between the block size and arithmetic intensity (or peak SRAM usage), allowing the programmer to maximize utilization of compute engines while respecting the memory constraint.
We summarize the major contributions as follows.
  • We build a high-performance matmul, NeuronMM, for LLM inference on Trainium. NeuronMM is open-sourced and adds a key milestone to the Trainium eco-system.
  • We introduce a series of techniques customized to Trainium to reduce data movement across the software-managed memory hierarchy, maximize the utilization of SRAM and compute engines, and avoid expensive matrix transpose.
[…]
2. Background
2.1. SVD for Weight Compression
SVD is a well-established technique for approximating high-dimensional matrices with low-rank representations (demmel1997applied). Given a weight matrix $W$, SVD factorizes it into three matrices: $U$, $\Sigma$, and $V$, such that $W=U\Sigma V^{\top}$. By retaining only the top-$k$ singular values in $\Sigma$ and their corresponding singular vectors in $U$ and $V$, one obtains a low-rank approximation $W\approx U_{k}\Sigma_{k}V_{k}^{\top}$. This approximation preserves the most informative components of $W$ while substantially reducing the number of parameters to represent $W$. As a result, SVD is particularly well-suited for compressing the large weight matrices in the linear layers of LLMs, where parameter reduction directly improves efficiency with tolerable accuracy loss.
Applying SVD involves two steps: matrix factorization and fine-tuning. We perform SVD offline on invariant weight matrices, and the resulting low-rank factors are used during inference for efficient matmuls. This design specifically targets LLM inference, where weights remain fixed and compression directly improves performance. In contrast, LLM training continuously updates weights, making SVD-based factorization impractical. Hence, NeuronMM is optimized specifically for inference.
2.2. AWS Trainium
AWS Trainium is a custom silicon chip designed to accelerate deep learning workloads. It adopts a systolic array–based architecture with rich hardware heterogeneity. Each Trainium chip integrates two NeuronCores. Each NeuronCore functions as an independent heterogeneous compute unit composed of a rich set of specialized engines designed for different operations, such as tensor engine, scalar engine, vector engine, and GPSIMD engine. In addition, Trainium includes DMA engines that can transfer data between HBM and on-chip SRAM in Trainium. Those engines operate in parallel, enabling Trainium to efficiently support diverse deep learning tasks (aws_trainium_arch). In addition, we particularly focus on the tensor engine for matmul in the following discussion.
We use Neuron Kernel Interface (NKI) (aws2024neuron), a bare-metal language and compiler for directly programming NeuronDevices available on AWS Trainium, in our study.
Tensor engine. It accelerates matmuls by reading input tiles from SBUF (an on-chip SRAM on Trainium) and writing output tiles to PSUM (another on-chip SRAM on Trainium). The tensor engine is organized as a $128\times 128$ systolic array of processing elements, defining a partition dimension ($P=128$), where each partition maps to a memory partition in SBUF or PSUM. To fully exploit parallelism across the 128 processing units, the contraction dimension of a matmul must align with the partition dimension, allowing each partition to process a distinct tile of data concurrently.
Each tile-level matmul involves two input matrices, named as the stationary (left-hand side) matrix and moving (right-hand side) matrix (using AWS terms). The tensor engine loads the stationary matrix into its internal storage and streams the moving matrix across it. Due to the hardware’s systolic array design, the stationary matrix must be consumed in a transposed layout (aws_trainium_arch). For clarity, we denote the low-level instruction nki.isa.nc_matmul (aws_nki_matrix_multiplication) as NKIMatmul, which carries out this tile-level operation. Thus, computing the product of $AB$ is formulated as NKIMatmul(stationary = $A^{T}$, moving = $B$). The stationary matrix, which remains fixed during the computation, is transposed so that its rows align with the columns of the moving matrix, ensuring both input tiles share the same first dimension which corresponds to the partition dimension in SBUF.
Figure 1. NeuronCore memory hierarchy on Trainium with bandwidth and memory size.
Memory heterogeneity. A NeuronCore is associated with three types of memory (Figure 1): 16 GB off-chip High Bandwidth Memory (HBM), 24 MB on-chip State Buffer (SBUF), and 2 MB on-chip Partial Sum Buffer (PSUM). SBUF serves as the primary on-chip data buffer while PSUM serves as the dedicated accumulation buffer for the tensor engine. Both SBUF and PSUM are two-dimensional, each consisting of 128 partitions. Computation proceeds by loading data from HBM into SBUF, where the data is accessible by all engines. Once the computation is completed, the final results are written back to HBM. This explicit programming-model shifts responsibility to software: efficient tiling and placement are essential to exploit on-chip locality. Inefficient management instead leads to excessive HBM traffic, longer DMA transfers, and tensor engine stalls, directly limiting Trainium’s performance.
[…]
Figure 2. Matmul tiling on Trainium (mathematical view).
Figure 3. Matmul tiling on Trainium (hardware view).
Three-level hierarchical data layout. For clarity, given a matrix $A\in\mathbb{R}^{M\times N}$, we define a hierarchical three-level data layout, ordered from the finest to coarsest: the Trainium-native tile, the logical block, and the matrix-spanning strip.
The tile is the fundamental, hardware-native unit of data processed by a single NKIMatmul instruction. The tensor engine imposes strict, hardware-defined maximum dimensions on these tiles, and these constraints differ for the two matrix operands (aws_trainium_arch). Specifically, a tile for the stationary matrix cannot exceed dimensions of (128, 128) for its partition and free axes, respectively. In contrast, a tile for the moving matrix can have a larger free axis, with its dimensions limited to (128, 512). When a matrix is larger than these hardware limits, it must be divided into multiple tiles for processing.
The block is a higher-level, logical software construct composed of one or more computational tiles. Specifically, a block groups a $t_{M}\times t_{N}$ grid of tiles, where $t_{M}$ and $t_{N}$ denote the number of tiles along two dimensions. This forms a block with dimensions $(B_{M},B_{N})$, where $B_{M}=t_{M}\times T_{M}$, $B_{N}=t_{N}\times T_{N}$, and $T_{M},T_{N}$ is the tile size of Trainium. When a matrix $A$ is partitioned into blocks, we denote the block at the $i^{\text{th}}$ row and $j^{\text{th}}$ column as $A_{ij}$. We use lowercase notation to denote individual tiles; for example, $a_{kl}$ refers to the tile at the $k^{\text{th}}$ row and $l^{\text{th}}$ column of $A$.
The strip is a collection of blocks that span one entire dimension of $A$. A row strip, corresponding to the $i^{\text{th}}$ row of blocks in matrix A, is denoted as $A_{i*}$. Similarly, a column strip can be denoted as $A_{*j}$. A strip is thus a set of blocks (e.g., $A_{i*}=\{A_{i1},A_{i2},\dots\}$) that form a sub-matrix, such as the one with a shape $(B_{M},N)$ for a row strip in a matrix with the dimension $(M,N)$.
The block-level multiplication in Trainium repeatedly performs the following three steps: (1) loading a pair of tiles from HBM to SBUF (one of the tiles must be transposed), (2) multipling the two tiles, and (3) storing the multiplication result into PSUM. PSUM is used to accumulate the tile-multiplication results from all tiles in the input blocks. We use MatMulBlock($s,m$) to denote this block-level matmul, where the stationary matrix $s$ multiplies the moving matrix $m$ on SBUF, and the result is accumulated into PSUM.
Matmul tiling on Trainium. PSUM serves as a dedicated landing buffer for the tensor engine, with near-memory accumulation capabilities that enables read-accumulate-write operations at a fine granularity of every 4B memory element. The accumulation mode of PSUM is particularly useful for large matmul, especially when the matrices have a high contraction dimension (usually the inner dimension in a dot product). For instance, consider a matmul operation where the input tensors have dimensions: x.shape = [128, 512] and y.shape = [512, 512]. As illustrated in Figures 2, the input matrix can be partitioned through tiling, resulting in sliced input tiles [x0, x1, x2, x3] and [y0, y1, y2, y3]. The final output is obtained by computing individual partial sums, such as output0 = matmul(x0, y0), followed by the accumulation of these intermediate outputs: output = output0 + … + output3.
Figure 3 shows that PSUM efficiently supports both intra-tile matmul and inter-tile result accumulation. Specifically, the first tensor-engine instruction writes its output directly into a designated PSUM bank, while subsequent instructions incrementally accumulate their results onto the existing content in the same PSUM bank. With eight PSUM banks allocated per partition, the tensor engine is capable of maintaining up to eight independent matmul accumulation groups concurrently. This architecture not only enhances the flexibility of matmul instruction scheduling on the tensor engine, but also enables overlap between the tensor engine operations and computation on other engines.
3. Motivation
Enabling high-performance matmul on Trainium faces multiple challenges. We discuss them in this section to motivate our work.
3.1. Challenge 1: I/O Bottleneck
As established in Section 2.1, SVD transforms a matmul $XW$ into a three-matrix chain $XUV$. However, executing this new formulation sequentially on Trainium introduces serious inefficiencies if not co-designed with the architecture.
Timeline
Figure 4. The Neuron Profiler view of up_projection matmul in Deepseek-V3 with SVD-compression. Directly computing on the SVD-compressed weight matrices sequentially leads to low SBUF and PSUM utilization and reduced Model Float Utilization (MFU). Frequent idle periods in the MFU indicate that the tensor engine is underutilized while waiting for data transfers and data preparation to complete.
Executing the SVD-compressed $XUV$ computation with a standard, sequential approach on Trainium results in poor hardware utilization. Figure 4 shows a Neuron Profiler trace of the SVD-compressed up_projection matmul in Deepseek-V3 (liu2024deepseek). For an input length of 4096, with hidden size 7168 and intermediate size 18432, the up_projection matrix ([7168, 18432]) is factorized into two low-rank matrices: $U$ ([7168, 4096]) and $V$ ([4096, 18432]). This transforms the original matmul ([4096, 7168] $\times$ [7168, 18432]) into a chain of three multiplications: [4096, 7168] $\times$ [7168, 4096] $\times$ [4096, 18432]. As highlighted in Figure 4, this decomposition results in frequent idle gaps in Model FLOPs Utilization (MFU) and inefficient use of the on-chip SBUF, indicating that the tensor engine is often stalled waiting for data.
The root cause is the materialization of the intermediate result. Sequential execution of $XUV$ first computes the intermediate matrix $Y=XU$, writes it from the fast on-chip SBUF to the slower HBM, and then reloads it into SBUF for the second multiplication $O=YV$. This load-compute-store cycle creates a severe I/O bottleneck inherent to Trainium’s systolic-style architecture. Compared to the original up_projection, the SVD-compressed version increases direct memory access (DMA) transfer time by 65% and more than doubles the traffic between HBM and SBUF.
1
2
Inputs: Matrices $X^{T}\in\mathbb{R}^{K\times M}$, $U\in\mathbb{R}^{K\times r}$, $V\in\mathbb{R}^{r\times N}$ in HBM.
Outputs: Result Matrix $O\in\mathbb{R}^{M\times N}$
3 Function MatMulBlock($A,B$):
  // Calculate block matrix multiplication
4  return $A\times B$;
5
6
7Initialize output $O$ in HBM;
8
9for $m\leftarrow 1$ to $\lceil M/B_{M}\rceil$ do
10  for $n\leftarrow 1$ to $\lceil N/B_{N}\rceil$ do
11     $O_{mn}\leftarrow\mathbf{0}$;
12     for $p\leftarrow 1$ to $\lceil r/B_{r}\rceil$ do
13        $Y_{mp}\leftarrow\mathbf{0}$;
14        for $k\leftarrow 1$ to $\lceil K/B_{K}\rceil$ do
15           Load blocks $X^{T}_{mk}$ and $U_{kp}$ from HBM;
16           $Y_{mp}\leftarrow Y_{mp}+\textnormal{{MatMulBlock}}(U_{kp},X^{T}_{mk})$;
17
18       $O_{mn}\leftarrow O_{mn}+\textnormal{{MatMulBlock}}(Y_{mp},V_{pn})$;
19
20    Write $O_{mn}$ back to $O$ on HBM;
21
22
23return $O$;
Algorithm 1 Naive kernel fusion for $Y=XUV$.
3.2. Challenge 2: Recomputation
To implement $XUV$ with minimum data movement, we fuse the two matmuls into a single kernel. We call our approach the naïve fused kernel. Instead of materializing the intermediate result $Y=XU$ in HBM, the kernel recomputes $Y$ on the fly for each block of the final output. While this eliminates intermediate HBM writes, it replaces the I/O overhead with a penalty in computation and HBM reads.
Algorithm 1 details our naive fused kernel. The computational penalty arises because the calculation of the intermediate block $Y_{mp}$ (lines 9–11) is nested inside the main loop over $n$ (lines 5-13). As a result, the kernel redundantly recomputes the same $Y_{mp}$ for every output block $O_{mn}$ in a row. The computation of each output block $O_{mn}$ can be expressed as the following nested summation:
(1) $\small O_{mn}=\sum_{p=1}^{\text{NumBlocks}_{r}}\left(\underbrace{\left(\sum_{k=1}^{\text{NumBlocks}_{K}}X_{mk}U_{kp}\right)}_{\text{Intermediate block }(XU)_{mp}}V_{pn}\right)$
This formula reveals the source of recomputation. The inner summation, which calculates each block of the intermediate matrix $(XU)_{mp}$, depends only on the row-block index $m$ and the reduction-block index $p$; it is independent of the output column-block index $n$. Yet, the naive fused kernel re-calculates this inner sum for every output block in the row strip $O_{m,*}$ ($O_{m,1},\dots,O_{m,\text{NumBlocks}_{N}}$). As a result, the same intermediate blocks are recomputed unnecessarily, incurring a penalty factor of $\text{NumBlocks}_{N}=\lceil N/B_{N}\rceil$.
This recomputation arises solely from Trainium’s on-chip memory limits. Avoiding recomputation would require setting $B_{N}=N$, producing an entire row-block $(B_{M},N)$ at once. But for LLM-scale matrices (e.g., $N=16384$), such a block cannot fit within the 24 MB SBUF. In practice, we are therefore constrained to choose a much smaller $B_{N}$ (e.g., 512), forcing repeated recomputation. Worse, each recomputation triggers additional I/O: source blocks of $X$ and $U$ are repeatedly reloaded from HBM, inflating memory traffic and erasing the benefits of kernel fusion.
[…]
Table 1. Evaluation of the sequential matmul and naïve kernel fusion for matmul.
Metric Sequential Matmul Naive Kernel fusion
Total Time (ms) 1.57 18.06
Model FLOPs (GFLOPs) 85.90 343.60
Memory Footprint (MB) 298.66 3140.42
3.3. Challenge 3: Transpose Overhead
Transpose overhead on Trainium stems from the systolic-array design of its tensor engine, which requires the stationary matrix in a matmul to be supplied in a transposed layout. This requirement conflicts with the natural data flow of LLMs and introduces two types of overhead: (1) I/O transposes on tensors entering or leaving a kernel, and (2) intermediate transposes on temporary results produced on-chip. Our goal is to eliminate intermediate transposes entirely and minimize the impact of I/O transposes.
I/O transpose. An I/O transpose occurs when a tensor’s layout expected by an NKI compute API differs from the layout stored in HBM which is constrained by the surrounding LLM computation graph. In the $XUV$ operator, for example, the input activation matrix $X$ must remain in its natural, non-transposed layout to integrate with the LLM data flow. However, for the first matmul $Y=XU$, the tensor engine requires the stationary input to be $X^{T}$. Thus, transposing $X$ is unavoidable, making it an I/O transpose.
Intermediate transpose. An intermediate transpose occurs when the output of one NKI kernel must be reshaped before it can be consumed by another kernel due to a layout mismatch. In the sequential execution of $XUV$, the first matmul produces $Y=XU$, which serves as the stationary input for the second multiplication $O=YV$. Because the tensor engine requires $Y^{T}$, a naïve implementation must explicitly transpose $Y$ on chip. This extra step is costly, as it introduces additional memory movement and synchronization overhead.
(a) Block-aligned SVD
(b) TrainiumFusion
Figure 5. The overview of NeuronMM. (a) Block-aligned SVD. The weight parameters of the attention layers remain unchanged, while only the large matrices $W$ in the MLP layers are compressed using SVD. (b) TrainiumFusion. The weight $W$ is decomposed into $U$ and $V$, and the original matmul XWXW turns into XUVXUV. The kernel leverages caching, implicit transposition, and blocking to enable efficient matmul, thereby reducing data movement between off-chip HBM and on-chip SRAM (SBUF and PSUM).
4. Design
Figure 5 provides an overview of NeuronMM, a framework for accelerating LLMs on AWS Trainium. Table 2 summarizes the notations used in this section.
NeuronMM compresses the large weight matrices in MLP layers using block-aligned SVD and restores accuracy through Low-Rank Adaptation (LoRA) (hu2022lora) fine-tuning. It further provides high-performance NKI kernels, termed TrainiumFusion, that execute matmuls on compressed models efficiently by exploiting Trainium’s architectural features.
Table 2. Notations.
Notation Description
Matrix Operations & Dimensions
$W\approx UV$ The weight matrix $W$ is approximated by the product of two low-rank matrices.
$X\in\mathbb{R}^{M\times K}$ The input activation matrix.
$U\in\mathbb{R}^{K\times r},V\in\mathbb{R}^{r\times N}$ The low-rank matrices from SVD.
$M$ The input sequence length.
$K,N$ The hidden size and intermediate size in MLP layer of LLMs.
$r$ The rank of the SVD-decomposed matrices.
Data Layout Hierarchy
$B_{M},B_{K},B_{r},B_{N}$ The sizes of a block along dimensions.
$T_{M},T_{N}$ The sizes of a tile along dimensions.
$A_{ij}$ The block at the $i$-th row and $j$-th column of matrix A.
$A_{i*}$ The row strip composed of all blocks in the $i$-th row of matrix A.
$A_{*j}$ The column strip composed of all blocks in the $j$-th column of matrix A.
4.1. Block-Aligned SVD
Leveraging existing work. Following the standard workflow of post-training LLM compression methods (yuan2023asvd; wang2024svd; ding2025dipsvd), the weight matrix $W$ is first scaled by a matrix $S$ to capture the influence of input activations. $S$ is derived from a random set of input sentences. Specifically, for each MLP layer in LLMs, we record input activations $X$ using forward hooks, compute the covariance matrix, and apply Cholesky decomposition:
(2) $SS^{\top}$ $=\text{Cholesky}(X^{\top}X)$
where $S$ is a lower triangular matrix with all positive diagonal elements. NeuronMM then performs SVD on $WS$ and reconstructs $W$ using $S^{-1}$.
(3) $W=(WS)\cdot S^{-1}=(U\Sigma V^{\top})\cdot S^{-1}\approx U_{r}\Sigma_{r}V_{r}^{\top}\cdot S^{-1}$
where $W\in\mathbb{R}^{k\times n}$, $U_{r}^{\prime}\in\mathbb{R}^{k\times r}$, $V_{r}^{\prime}\in\mathbb{R}^{r\times n}$ and $r$ denotes the top-r singular values. The choice of $r$ is crucial for balancing model accuracy and compression ratio.
We define the compression ratio as follows.
(4) $\text{ratio}=1-\frac{\text{Model}\_\text{size}_{\text{compressed}}}{\text{Model}\_\text{size}_{\text{original}}}$
where $\text{Model}\_\text{size}_{\text{compressed}}$ and $\text{Model}\_\text{size}_{\text{original}}$ denote the model parameter counts with and without compression. This definition follows the existing work on LLM compression (wang2024svd; wang2025svd; li2025adasvd; yuan2023asvd; ding2025dipsvd).
Block alignment. Unlike prior work, we introduce a block-aligned rank selection strategy that maximizes tensor engine utilization by coupling the compression ratio with the tensor engine’s $tile_{s}ize$. The rank $r$ is computed as follows.
(5) $r=\left\lfloor\frac{k\times n\times(1-ratio)}{(k+n)\times block\_size}+\alpha\right\rfloor\times block\_size$
where $k$ and $n$ are the dimensions of $W$, and block_sizeblock\_size is an integer multiple of tile_sizetile\_size. $\alpha$ is a rounding threshold that adjusts the required number of blocks, with $\alpha=0.5$ corresponding to standard rounding. This formulation ensures that the choice of $r$ both satisfies the target compression ratio and aligns with hardware tile boundaries, avoiding intra-block padding and improving utilization of SBUF and the tensor engine. With $r$ determined,   $U_{r},\Sigma_{r},V_{r}^{\top}$, and $S^{-1}$ are consolidated into $U_{r}^{\prime}$ and $V_{r}^{\prime}$.
(6) $\footnotesize U_{r}^{\prime}$ $=U_{r}\cdot\sqrt{\Sigma_{r}}$
(7) $V_{r}^{\prime}$ $=\sqrt{\Sigma_{r}}\cdot V_{r}^{\top}\cdot S^{-1}$
(8) $W$ $\approx U_{r}^{\prime}V_{r}^{\prime}$
This transformation represents each weight matrix as the product of two low-rank matrices, reducing parameters from $kn$ to $r\times(k+n)$.
Although SVD substantially reduces the size of the weight matrix, it inevitably introduces accuracy loss. To recover accuracy, we apply LoRA fine-tuning to the compressed weights, similar to the prior work (wang2024svd; li2025adasvd). During fine-tuning, we freeze the compressed weights $U_{r}^{\prime}$ and $V_{r}^{\prime}$, and fine-tune them with LoRA:
(9) $U\leftarrow U_{r}^{\prime}+B_{u}A_{u},\quad V\leftarrow V_{r}^{\prime}+B_{v}A_{v}$
where $A_{u},B_{u},A_{v}$, and $B_{v}$ are the trainable matrices used to adapt the model via LoRA. After fine-tuning, we incorporate the matrices $B_{u}A_{u}$ and $B_{v}A_{v}$ into $U_{r}^{\prime}$ and $V_{r}^{\prime}$, respectively, to form the final compressed weight matrices $U$ and $V$.
4.2. TrainiumFusion
After SVD, we build a $XUV$ NKI kernel using new kernel fusion techniques in Trainium.
4.2.1. XUV NKI Kernel
We introduce three techniques: caching, implicit transposition, and blocking, to overcome the challenges of I/O bottlenecks and recomputation. The main idea is to fuse the $XUV$ chain into a two-stage computation that executes entirely within the on-chip SBUF. First, we compute a strip of the intermediate product, using implicit transposition by reordering the inputs to the NKIMatmul primitive to directly generate its transpose, $(XU)^{T}$. This on-chip result is then immediately consumed in the second stage, where it is multiplied with a corresponding strip of $V$ to produce a block of the final output. This fused dataflow avoids intermediate data transfer between HBM and SBUF, eliminates the recomputation penalty, and removes the intermediate transpose, as illustrated in Figure 5(b). We give more details as follows.
Caching. NeuronMM uses a capacity-aware caching strategy to eliminate recomputation penalty. The kernel calculates an entire row strip of the intermediate matrix, $(XU)_{m*}$, and caches it in a dedicated buffer within the on-chip SBUF. This cached strip is then efficiently reused for the subsequent multiplications with all corresponding column strips of the $V$ matrix ($V_{*1},V_{*2},\dots$) to produce every output block ($O_{m1},O_{m2},\dots$) to eliminate the recomputation. This on-chip caching is feasible within the SVD-based LLM, because the intermediate strip’s memory usage is manageable. In particular, its shape is $(B_{M},r)$, where $B_{M}$ is a block size (e.g., $1024$) and $r$ is the SVD rank. The resulting buffer size (e.g., $1024\times r\times 2$ bytes for float16 tensor) fits within the 24MB SBUF, leaving enough space for other necessary data blocks.
[…]
Blocking. We compute $XUV$ by blocks as shown in Figure 5(b). The dataflow is structured around a main outer loop that processes the input matrix $X$ one row strip at a time. Within each iteration of this outer loop, the computation proceeds in two phases within inner loops. First, the kernel computes and caches the entire intermediate strip, $(XU)_{m*}^{T}$, in SBUF. To do this, its inner loops load corresponding blocks of $X$ and $U$ from HBM. The $X$ block is transposed in transit by the DMA engine, and the blocks are multiplied, with the result $(XU)_{mp}^{T}$ accumulated in PSUM before being stored in the SBUF cache. In the second phase, another set of inner loops iterates through the blocks of matrix $V$, loading them from HBM and multiplying them with the pre-computed blocks fetched from the cached strip in SBUF. The final result for the output block, $O_{mn}$, is accumulated in PSUM and then written back to HBM.
The block size can impact kernel performance significantly because there is a trade-off between computational efficiency and on-chip memory usage. We model the trade-off with two metrics – Arithmetic Intensity and Peak SBUF Usage. Assume that we have the inputs $X\in\mathbb{R}^{M\times K}$, $U\in\mathbb{R}^{K\times r}$ and $V\in\mathbb{R}^{r\times N}$ in HBM, and $s$ is the size of the data type of inputs. The arithmetic intensity (Equation  10) is defined as the ratio of total FLOPs to HBM traffic. The HBM traffic is composed of initial reads of $X$, final writes of $O$, and repeated reads of $U$ and $V$ for each of the $M/B_{M}$ row strips.
Arithmetic Intensity $=\frac{2Mr(K+N)}{s\cdot\left(M(K+N)+\frac{M}{B_{M}}r(K+N)\right)}$
(10) $=\frac{2r}{\left(1+\frac{r}{B_{M}}\right)\cdot s}$
The peak SBUF usage is determined by the maximum memory required across the kernel’s two computational phases: first computing the intermediate strip $(XU)_{m*}$, and second, using that strip to produce the final output blocks. The peak requirement is the maximum of the SBUF footprints in these two phases, formulated as follows.
Peak SBUF Usage $=\max\left((B_{M}r+B_{M}B_{K}+B_{K}B_{r}),\right$.
$\quad\left.(B_{M}r+B_{r}B_{N}+B_{M}B_{N})\right)\cdot s$
(11) $=(B_{M}r+(B_{M}+B_{r})\cdot\max(B_{K},B_{N}))\cdot s$
Equations 10-11 model a trade-off: the arithmetic intensity increases with a larger $B_{M}$ due to better data reuse, while the peak SBUF usage also grows as it must hold larger blocks. Therefore, an optimal block size must be large enough to maximize arithmetic intensity and saturate the tensor engine, yet small enough to fit within the 24 MB SBUF capacity to avoid memory spills. As shown in Section LABEL:sec:block-exp, our experimental results validate the two models.
Using Equations 10-11, we determine the optimal block size. By the roofline model, a kernel becomes compute-bound when its arithmetic intensity exceeds a hardware-specific threshold (williams2009roofline)—for a Trainium NeuronCore with bfloat16 data, this threshold is 222 Flops/Byte (nki_mm2025). Setting the arithmetic intensity (Equation 10) to this threshold yields the minimum block size $B_{M}$ required to saturate the tensor engine. Starting with this $B_{M}$, the peak SBUF usage model (Equation 11) can then identify block combinations that maximize data reuse within the 24 MB SBUF capacity.
4.2.2. Kernel Input and Output Layout
Our $XUV$ NKI kernel must integrate seamlessly into the LLM in a high-level framework like PyTorch or vLLM, where tensors typically maintain a standard, non-transposed layout. This presents a challenge, as Trainium’s systolic array architecture requires the stationary matrix in a multiplication to be transposed. To manage this layout mismatch, our strategy is twofold: we aim to completely eliminate transposes between kernels (intermediate transposes) and minimize the performance impact of unavoidable transposes at the kernel’s boundary (I/O transposes).
Input layout. The unavoidable I/O transpose of the input matrix $X$ can be addressed in two ways. One option is to load data at full DMA bandwidth and perform the transpose on the tensor engine, but this wastes valuable compute cycles. The alternative is to let the DMA engine transpose data on-the-fly, which lowers effective DMA bandwidth but frees the tensor engine for matmul. We adopt the latter approach: the DMA engine handles transposes, while the tensor engine remains dedicated to computation. This strategy overlaps communication with computation, as the DMA engine can load and transpose the next tile while the tensor engine processes the current one.
Output layout. The kernel’s output layout is chosen to eliminate downstream transpose operations, depending on the output’s consumer. If the output is to be consumed by another NKI kernel that requires a transposed, stationary matrix, we produce $O^{T}$ directly. If the output is passed back to the high-level model, which expects a standard tensor layout, we produce the non-transposed output $O$. By swapping the order of the stationary and moving matrices in NKIMatmul, we can produce either $O$ or $O^{T}$ with no performance overhead.
4.2.3. MLP NKI Kernel
We introduce a specialized NKI kernel for SVD-compressed MLP layers, which extends the $XUV$ kernel design to a multi-stage operation.
The computation in an MLP layer, such as one using SwiGLU, involves three matmuls. First, two parallel linear transformations—a “gate” projection and an ”up” projection—are applied to the input tensor $X$. The gate’s output is passed through a SiLU activation function and then combined with the up-projection’s output via an element-wise multiplication. This intermediate result is then passed through a “down” projection to produce the MLP layer’s output.
Our implementation (Algorithm 3) maps the MLP computation onto two NKI kernels (lines 6-7) derived from the $XUV$ kernel. The first stage, UpGateProjection kernel (Algorithm 2) extends the XUV kernel to compute the “gate” and “up” projections in parallel (lines 6-20). It then uses Trainium’s Scalar and Vector Engines to perform $\texttt{SiLU}(\text{gate})\odot\text{up}$ on-chip SBUF and writes the transposed result to HBM (line 21). The second stage, DownProjection, uses the $XUV$ kernel that accepts a pre-transposed input and produces a standard-layout output. This stage seamlessly consumes the transposed output from the first stage without intermediate transpose, and writes the final result to HBM in the standard layout required by subsequent LLM layers.
4.3. Discussions
Focus on MLP in LLM. A typical transformer layer in LLM consists of the attention and MLP, both of which employ matmul. We apply NeuronMM to MLP because of the following two reasons. First, the parameters in MLP account for the majority of parameters in LLM. For example, in Llama-3.1-8B, the parameters in MLP takes 70% of the overall parameters. Hence, working on MLP can bring larger reduction in inference time and LLM size. Second, applying NeuronMM to attention and MLP, we find big loss in LLM accuracy, even though we go through rigorous fine-tuning process to restore accuracy. Hence, we apply NeuronMM to MLP alone.
1
Inputs: Matrices $X\in\mathbb{R}^{M\times K}$, $U^{\text{gate}},U^{\text{up}}\in\mathbb{R}^{K\times r}$, $V^{\text{gate}},V^{\text{up}}\in\mathbb{R}^{r\times N}$ on HBM.
Outputs: Matrix $Y^{T}\in\mathbb{R}^{N\times M}$.
Note: Block matrix multiplication $A\cdot B$ refers to the hardware instruction MatmulBlock(s=A,m=B)(s=A,m=B).
2
3Function UpGateProjection($X,U^{\text{gate}},V^{\text{gate}},U^{\text{up}},V^{\text{up}}$):
4  Allocate $Y^{T}\in\mathbb{R}^{N\times M}$ on HBM;
5  for $m\leftarrow 1$ to $\lceil M/B_{M}\rceil$ do
6     Allocate $G_{m*},U_{m*}\in\mathbb{R}^{B_{M}\times r}$ on SBUF;
7     for $p\leftarrow 1$ to $\lceil r/B_{r}\rceil$ do
8        Initialize $G^{T}_{mp},U^{T}_{mp}\in\mathbb{R}^{B_{r}\times B_{M}}$ on PSUM with $\mathbf{0}$;
9        for $k\leftarrow 1$ to $\lceil K/B_{K}\rceil$ do
10           Load blocks $X^{T}_{mk}$, $U^{\text{gate}}_{kp}$ and $U^{\text{up}}_{kp}$ from HBM;
11
[…]
13
[…]
15
16       Write $G^{T}_{mp},U^{T}_{mp}$ to the $p$-th block of $G_{m*},U_{m*}$;
17
18    for $n\leftarrow 1$ to $\lceil N/B_{N}\rceil$ do
19        Initialize $G^{T}_{mn},U^{T}_{mn}\in\mathbb{R}^{B_{N}\times B_{M}}$ on PSUM with $\mathbf{0}$;
20        for $p\leftarrow 1$ to $\lceil r/B_{r}\rceil$ do
21           Load blocks $V^{\text{gate}}_{pn}$ and $V^{\text{up}}_{pn}$ from HBM;
22
23          1exFetch cached block $G^{T}_{mp},U^{T}_{mp}$ from $G_{m*},U_{m*}$;
24
[…]
26
[…]
28
29       Write $Y_{mn}^{T}\leftarrow\texttt{SiLU}(G_{mn}^{T})\odot U_{mn}^{T}$ to HBM;
30
31
32 return $Y^{T}$
Algorithm 2 MLP up-projection kernel
1
2Require: Matrices $X\in\mathbb{R}^{M\times K}$, $U^{\text{gate}},U^{\text{up}}\in\mathbb{R}^{K\times r}$, $V^{\text{gate}},V^{\text{up}}\in\mathbb{R}^{r\times N}$, $U^{\text{down}}\in\mathbb{R}^{N\times r_{2}}$, $V^{\text{down}}\in\mathbb{R}^{r_{2}\times K}$ on HBM. Return: Matrix $Z\in\mathbb{R}^{M\times K}$.
3 Note: Boolean is_XT and req_OT: input $X$ is pre-transposed, required output is $O^{T}$.
4 Function XUV_Kernel($X,U,V,\text{is\_XT},\text{req\_OT}$):
5  return XUVXUV or $(XUV)^{T}$
6
7Function FusedMLP(X, $U^{\text{gate}},U^{\text{up}},V^{\text{gate}},V^{\text{up}},U^{\text{down}},V^{\text{down}}$):
8
[…]
Algorithm 3 SVD-compressed MLP layer
5. Evaluation
5.1. Experimental Setup
Implementation. We develop NeuronMM on top of NeuronX Distributed Inference library (aws-neuron-neuronx-distributed-inference) and implement MLP kernels based on NKI (nki2025). We evaluate NeuronMM on an trn1.2xlarge instance of Amazon Elastic Compute Cloud (Amazon EC2) equipped with AWS Trainium accelerators, running the Deep Learning Amazon Machine Images (AMI) Neuron (Ubuntu 22.04). We use a single NeuronCore with 16GB HBM on a Trainium chip for evaluation, because using more than one core for inference based on tensor/pipeline/ data parallelism is not fully supported in NKI yet.
Models and datasets. We test LLMs that fit entirely into the HBM and are currently supported by NeuronX Distributed Inference library, including Llama-3.2-1B, Llama-3.2-3B, Qwen3-1.7B, and Qwen3-4B. We evaluate NeuronMM with nine datasets, covering three language modeling datasets (WikiText-2 (merity2016pointer), PTB (marcinkiewicz1994building), and C4 (raffel2020exploring)) and six common sense reasoning datasets (OpenBookQA (mihaylov2018can), WinoGrande (sakaguchi2021winogrande), PIQA (bisk2020piqa), HellaSwag (zellers2019hellaswag), ARC-e, and ARC-c (clark2018think)). For fine-tuning with LoRA, we use the yahma/alpaca-cleaned (yahma_alpaca_cleaned).
5.2. Evaluation of $XUV$ Kernel
We compare NeuronMM against two baselines: NKI $XW$ and NKI $XUV$. Both baselines use the state-of-the-art matmul kernel implemented by AWS official (aws_nki_matrix_multiplication).
Table 3. Performance under different block size $B_{M}$.
$B_{M}$ 128 256 512 1024 2048 4096
Total Time (ms) 31.25 16.02 11.02 10.99 11.07 12.50
Arithmetic Intensity (flops/byte) 124.12 240.94 455.10 819.17 1280.50 512.95
SBUF Usage (%) 19.54 51.69 80.07 90.05 96.35 98.96
Spill Reload (MB) 0 0 0 0 29.19 931.00
Spill Save (MB) 0 0 0 0 10.53 266.00
5.3. Evaluation of MLP Kernel
We integrate NeuronMM into the MLP layer of LLMs and evaluate its perplexity, accuracy and speedup on end-to-end inference. For language modeling tasks (Wiki2, PTB, and C4), we report perplexity (PPL), which measures the model’s uncertainty in predicting the next token, with lower values indicating better predictions. For common-sense reasoning tasks (the other six tasks), we report task accuracy.
To assess the trade-off between inference speed and accuracy degradation, we adopt the Speedup Degradation Ratio $\gamma$ (he2024matters):
(12) $\footnotesize\gamma=\frac{\text{Avg}_{\text{full}}-\text{Avg}_{\text{method}}}{\text{Speedup}_{\text{method}}-\text{Speedup}_{\text{full}}}$
where $\text{Avg}_{\text{full}}$ and $\text{Avg}_{\text{method}}$ denote the average accuracy of the full model and the performance-optimization method, respectively. A lower $\gamma$ reflects greater efficiency, capturing how well a method retains accuracy for each unit of speedup.
5.3.1. Impact of Compression Ratio
We evaluate Llama-3.2-1B under compression ratios from 0.05 to 0.5 on the WikiText-2 language modeling task and the ARC Easy commonsense reasoning benchmark. As shown in Figure 6, higher compression ratios reduce the number of model parameters but lead to increased perplexity and decreased accuracy. This demonstrates the fundamental trade-off between model size and predictive performance. The model degradation is acceptable under mild compression. However, as the compression ratio increases performance drops sharply. In particular, between 0.4 and 0.5, PPL surges from 38.87 to 51.29, and accuracy drops from 46% to 40%, indicating that compression has significantly harm the model’s capabilities. To ensure that the accuracy degradation stays within an acceptable range (wang2024svd; wang2025svd; li2025adasvd; yuan2023asvd; wang2025dobi), we adopt compression ratios of 0.1 and 0.2 for the end-to-end evaluation (Sections 5.3.3-5.3.4).
Figure 6. Model degradation with increasing compression ratios.
5.3.2. Impact of LoRA Fine-Tuning
Figure 7 presents the accuracy of Qwen-3-1.7B under compression ratios of 0.1 and 0.2 across six commonsense reasoning benchmarks. We can observe that while compression reduces accuracy, LoRA fine-tuning effectively restores it, keeping degradation negligible at these low ratios.
Figure 7. Accuracy degradation and recovery of Qwen-3-1.7B under different compression ratios on six common-sense reasoning datasets.
Table 4. Evaluation of NeuronMM across four LLMs and nine datasets under compression ratios of 0.1 and 0.2.
Model Compr Ratio PPL (↓) Accuracy (↑) mAcc (↑) Avg. Speedup (↑) $\gamma$ (↓)
Wiki2 PTB C4 Openb. ARC_e ARC_c WinoG. HellaS. MathQA
Llama-3.2-1B 0 9.75 15.40 13.83 0.26 0.66 0.31 0.61 0.48 0.29 0.43 1.00×
0.10 15.64 22.80 22.72 0.20 0.59 0.28 0.55 0.41 0.26 0.38 1.21× 25.27%
0.20 18.87 27.24 26.71 0.18 0.56 0.26 0.54 0.39 0.25 0.36 1.63× 11.24%
Llama-3.2-3B 0 7.82 11.78 11.29 0.31 0.74 0.42 0.70 0.55 0.35 0.51 1.00×
0.10 11.58 15.61 17.11 0.24 0.65 0.34 0.64 0.47 0.28 0.44 1.88× 8.66%
0.20 15.13 18.69 20.80 0.23 0.62 0.29 0.60 0.43 0.27 0.41 2.49× 7.20%
Qwen-3-1.7B 0 16.68 28.88 22.80 0.28 0.72 0.40 0.61 0.46 0.38 0.47 1.00×
0.10 15.43 25.00 23.16 0.27 0.72 0.39 0.62 0.43 0.34 0.46 1.41× 3.24%
0.20 17.05 26.97 25.14 0.25 0.70 0.37 0.58 0.42 0.31 0.44 1.74× 4.72%
Qwen-3-4B 0 9.75 15.40 13.83 0.29 0.80 0.51 0.66 0.52 0.47 0.54 1.00×
0.10 12.18 18.98 19.05 0.31 0.78 0.47 0.66 0.50 0.42 0.52 1.28× 6.93%
0.20 14.05 21.09 21.38 0.30 0.75 0.43 0.64 0.48 0.37 0.49 1.67× 7.41%
5.3.3. Impact of Different LLMs
To demonstrate the generability of NeuronMM across various LLMs, we evaluate it on four LLMs (Llama-3.2-1B, Llama-3.2-3B, Qwen3-1.7B, and Qwen3-4B) across 9 datasets. We report mean accuracy (mAcc) and average end-to-end speedup, with results summarized in Table 4. Across all LLMs evaluated, the SVD-compressed models retain accuracy largely comparable to the original, with mAcc drop $\leq$ 0.10 in every case, a level of loss generally considered acceptable (wang2024svd; wang2025svd; li2025adasvd; yuan2023asvd; wang2025dobi).
[…]
5.3.4. Impact on Inference TTFT and TPOT
[…]
Table 5. Module sizes of Llama-3.2-1B with and without SVD compression at compression ratio of 0.2.
Without SVD Compression Ratio 0.2
Module Size Module Size
up_proj [2048, 8192] up_u_proj [2048, 1280]
up_v_proj [1280, 8192]
gate_proj [2048, 8192] gate_u_proj [2048, 1280]
gate_v_proj [1280, 8192]
down_proj [8192, 2048] down_u_proj [8192, 1280]
down_v_proj [1280, 2048]
Table 6. Performance of Llama-3.2-1B under different compression ratios with batch size= 1, ISL = 1024, OSL = 1024. Ratio means compression ratio.
Ratio E2E Latency (s) Throughput TTFT (ms) TPOT (ms)
0 41.22 49.69 82.22 39.65
0.05 33.76 60.67 74.18 32.38
0.10 30.45 67.25 71.41 29.15
0.15 24.90 82.25 63.07 23.74
0.20 22.14 92.52 61.20 20.41
6. Related Work
SVD and LLM. LLM can have billions of parameters, making inference on resource-constrained hardware challenging. Various model compression techniques have been proposed to reduce their latency and memory footprint (dettmers2023qloraefficientfinetuningquantized; Frantar2023OPTQAQ; wu2023understandingint4quantizationtransformer; frantar2023sparsegpt; ma2023llmprunerstructuralpruninglarge; guo2023compressostructuredpruningcollaborative; ashkboos2024slicegpt; DynaBERTHou; miles2022informationtheoreticrepresentationdistillation; yang2022understandingprivilegedfeaturesdistillation). Recent works have explored various SVD-based approaches for model compression. For example, FWSVD (hsu2022language) introduces a weighted low-rank factorization scheme, while ASVD (yuan2023asvd) develops an activation-aware SVD technique that exploits layer-wise activation patterns to enhance compression effectiveness. SVD-LLM (wang2024svd; wang2025svd) further incorporates truncation-aware data whitening and layer-specific parameter updates, assigning unique ratios to individual weights. In contrast, Dobi-SVD (wang2025dobi) establishes a bijection mapping, allowing the model to automatically learn the best truncation points among layers. However, those works do not consider the characteristics of accelerator hardware and the additional matmul overhead introduced by SVD (e.g., memory spilling) as NeuronMM.
Performance optimization on Trainium. Recent studies (xue2024ninjallmfastscalablecosteffective; biran2025adaptiveorchestrationlargescaleinference) have demonstrated that AWS Trainium is promising to accelerate generative AI workloads. HLAT (fan2024hlat) demonstrates that, through a series of system-level optimizations and training strategies, Trainium can reduce training costs to 60% of those on p4d GPU instances while maintaining comparable model quality to GPU-based baselines. Complementing this, a study (fu2024distributed) details NeuronX Distributed Training (NxDT), quantifies scaling and efficiency against contemporary GPU baselines, and elucidates runtime and compiler support critical for stable large-cluster operation.
While most existing work leverages Trainium primarily for large-scale pretraining, very few studies investigate its use for inference. Inference workloads on Trainium often involve substantial data movement between off-chip and on-chip memory, which can become a significant performance bottleneck. This motivates our work, which focuses on mitigating these data movement costs to unlock the full inference potential of Trainium.
7. Conclusions
[…]